A wire is formed into a circle having a diameter oI need help answering 18, 19, and 20. Thanks!
Answer #1

18)a)τ= μ*B = NIABsinθ = 1*5.00*π*0.09^2*3.10*10^-3*sin90 = 0.4*10^-3 N.m

b) The energy of the system is given by U = − μ*B = − μBcosθ. The range of U depends on the fact that cos ranges between +1 and − 1. Hence the range of the potential energy is – 1.18* 10 ^− 4 J to1 .18 *10^ − 4 J

19) Magnetic field for the solenoid , B= μ0NI/L

1*10^-4= (8.85*10^-12*I*1180)/(0.440)     => I = 4.2*10^3 A

c) Magnetic field for the solenoid , B= μ0NI/L

10 = (8.85*10^-12*55*N)/(1.0)      => N=2.05*10^10 turns